A 95-g aluminum calorimeter is filled with ice taken from the freezer at -8.50C. The calorimeter is then filled with 100 g of steam at a temperature of 1030C. Assume the aluminum calorimeter has the same initial temperature as the ice. The final situation is observed to be all water at 90.00C.

  1. What was the mass of the ice?
Answer #1

Mass of the aluminium calorimeter = m1 = 95 g = 0.095 kg

Mass of the ice = m2

Mass of the steam = m3 = 100 g = 0.1 kg

Initial temperature of ice and calorimeter = T1 = -8.5 oC

Initial temperature of steam = T2 = 103 oC

Final temperature = T3 = 90 oC

Melting point of ice = T4 = 0 oC

Boiling point of water = T5 = 100 oC

Specific heat of aluminium = Ca = 900 J/(kg.oC)

Specific heat of ice = Ci = 2090 J/(kg.oC)

Specific heat of water = Cw = 4186 J/(kg.oC)

Specific heat of steam = Cs = 2010 J/(kg.oC)

Latent heat of fusion of ice = Li = 334000 J/kg

Latent heat of vaporization of water = Lw = 2260000 J/kg

The heat lost by the ice and the calorimeter is equal to the heat gained by the steam.

m1Ca(T3 - T1) + m2Ci(T4 - T1) + m2Li + m2Cw(T3 - T4) = m3Cs(T2 - T5) + m3Lw + m3Cw(T5 - T3)

(0.095)(900)(90 - (-8.5)) + m2(2090)(0 - (-8.5)) + m2(334000) + m2(4186)(90 - 0) = (0.1)(2010)(103 - 100) + (0.1)(2260000) + (0.1)(4186)(100 - 90)

8421.75 + 728505m2 = 230789

m2 = 0.305 kg

Mass of the ice = 0.305 kg